OpenStudy (yacoub1993):

Consider the leading term of the polynomial function. What is the end behavior of the graph? -3x5 + 9x4 + 5x3 + 3 a. The leading term is -3x5. Since n is odd and a is negative the end behavior is up and up b. The leading term is -3x5. Since n is odd and a is negative the end behavior is down and down c. The leading term is -3x5. Since n is odd and a is negative the end behavior is up and down d. The leading term is -3x5. Since n is odd and a is negative the end behavior is down and up

OpenStudy (anonymous):

I don't know. Maybe accessdenied can help.

OpenStudy (yacoub1993):

@AccessDenied need ur help please

OpenStudy (accessdenied):

End behavior is determined in a polynomial by the highest degree term, or in other words the \(ax^n\) whose value of n is largest. Do you recall any statements about how polynomial degree and coefficients affect this?

OpenStudy (yacoub1993):

c

OpenStudy (accessdenied):

For example, think about these simpler functions: y = x, y = -x, y = x^2, and y = -x^2. Odd powers tend to have a different right-end behavior from right-end behavior. Even powers tend to have the same end behavior both directions because you are inevitably squaring any negative or positive value. \( x^{2n} = (x^2)^n \) --- C. appears to be correct, then, because the left-end behavior is getting larger positive, and the right-end is getting larger negative. Think about putting in large positive or negative numbers in the leading term; -3( -big number)^3 = + 3 (big numbers). The odd power conserves the sign.

OpenStudy (accessdenied):

idifferent right-end behavior from left-end behavior. **